Cavo Corporation expects an EBIT of $21,000 every year forever. The company currently has no debt, and its cost of equity is 12 percent. The corporate tax rate is 35 percent. I have everything but PART C.

c-1.

What will the value of the company be if it takes on debt equal to 50 percent of its levered value?

c-2.

What will the value of the company be if it takes on debt equal to 100 percent of its levered value?
Answer #1

Ans :

(a) Ru = Cost of Equity =12 %

EBIT = $ 21,000

T = Tax rate = 35 %

Unlevered firm value = Vu = EBIT (1-T) / Ru

Vu = $ 21,000 ( 1-0.35) / 0.12

= $ 113,750.

(c)

c1. When debt is 50% of its levered value,

VL = Vu + T D

   VL = value of company of levered value

   D = 0.5 * VL

   Vu = $ 113,750.

Insert all the data, in formula VL = Vu + TD

VL = Vu + T * 0.5 * VL

VL(1-0.5T) = Vu

VL = Vu / (1-0.5T)

= $ 113,750/ (1-0.5*0.35)

= 113,750 / 0.825

= $ 137,878.78

c2 , D = 1 * VL

VL = Vu / (1-1*T)

= $ 113,750 / (1-0.35)

= $ 175,000